# Establishing a Baseline for Gaze-driven Authentication Performance in VR: A Breadth-First Investigation on a Very Large Dataset

### v1

@@ 3.1 Dataset
Of the $9202$ participants in GazePro, $3673$ self-reported as male, $5376$ as female, and $153$ as neither male nor female. The participants’ ages ranged from $13$ to $88$ years (median=$33$, IQR=$18$). All participants had normal or corrected-to-normal vision. Data collection went through the appropriate IRB approval protocol and received appropriate consent from participants. Additional participant details can be found in the supplementary material.
Gaze data in GazePro were collected at $72$ Hz on a VR headset with ET signal quality and other general characteristics equivalent to a Meta Quest Pro [6, 40]. Each recording contains binocular estimates of the optical and visual axes. The visual axis is a linear transformation of the optical axis. The parameters of the linear model (essentially, the angle kappa between the optical and visual axes) are fit on a per-user basis via a short calibration procedure.
@@ 3.3 Model evaluation
The first scenario is verification, for which we try to answer the question, “Is this user who they claim to be?” A potential user claims to be a specific person (e.g., the owner of a VR headset), their gaze signature is captured, and an accept/reject decision must be made based on the similarity of the gaze signatures captured from the claimed participant at enrollment time (e.g., during account creation) and from the potential user at verification time. We measure performance in this scenario using equal error rate (EER) and FRR at $1$-in-$50000$ FAR (denoted $\text{FRR}_{0.002\%}$).
The second scenario is identification, for which we try to answer the question, “Who is that person specifically?” A user’s gaze signature is captured, the similarity scores are computed between that user’s gaze signature and all gaze signatures from the population of enrolled participants, and the participant with the highest similarity is determined to be the best match. It is possible to reject a user if the highest similarity score is below some predetermined threshold, in which case it is assumed the user is not part of the enrolled population; but in our case, we evaluated identification rates only on users who were part of the enrolled population. We measure performance in this scenario using Rank-1 identification rate (IR).
@@ 4 Results
We highlight the verification performance under various conditions in Table 1. Unless otherwise noted, results were computed during a random saccade (jumping dot) task with the full training set ($6747$ users) and the full testing set ($2455$ users), using a model trained for $100$ epochs with a minibatch size of $m=256$ samples, and using $20$ seconds of gaze data for both enrollment and verification. In this case, we have $2359$ genuine matches and $5707331$ impostor matches (note that not every user was recorded for one or both of the tasks of interest). We assigned each experiment a unique ID to enable easier reference. Additional experiments can be found in the supplementary material.
Baseline.
Lohr & Komogortsev [27] used measures of the left eye’s visual axis. We did the same in Experiment 1 and achieved an EER of $8.53$% and a $\text{FRR}_{0.002\%}$ of $99.2$%.
@@
Experiment 2, which used the visual axis from both the left and right eyes, can be compared against our monocular baseline (Experiment 1). This experiment achieved an EER of $3.73$% and a $\text{FRR}_{0.002\%}$ of $93.1$%. These results were substantially better than the monocular baseline, indicating that having access to gaze estimates from both eyes indeed improved authentication performance over having gaze estimates from a single eye.
@@
Experiment 3, which used both the optical and visual axes from both the left and right eyes, can be compared against our visual axis baseline (Experiment 2). This experiment achieved an EER of $0.08$% and a $\text{FRR}_{0.002\%}$ of $17.7$%. These results were substantially better than the visual axis baseline, which makes sense considering that the user-specific calibration parameters can effectively be inferred from the difference between the two axes.
@@
We found that increasing the number of training epochs from $100$ to $1000$ yielded a large improvement in performance; however, given the large size of the GazePro dataset, this required more than $3$ weeks to train on a single NVIDIA V100 GPU. Simultaneously increasing the minibatch size from $m=256$ to $m=1024$ further improved performance.
In Experiment 5, we achieved $0.04$% EER and $4.1$% $\text{FRR}_{0.002\%}$ when training for $759$ epochs (terminated early due to time constraints) with a minibatch size of $m=1024$, and when using both the optical and visual axes from both the left and right eyes. These results were substantially better than Experiment 3 which trained for $100$ epochs with a minibatch size of $m=256$. The genuine and impostor similarity score distributions and receiver operating characteristic (ROC) curve for this experiment are presented in Figure 3.
[…]
@@
ET signal quality over a large population of people varies for many reasons, including but not limited to eye color, eye shape, presence of corrective lenses, makeup, and device placement on a user’s head [17]. Spatial accuracy, which is defined as the error between a user’s actual gaze and the gaze vector estimated by a gaze estimation pipeline, is one of the most important ET signal quality characteristics.
The subset of $8615$ participants who were recorded for both random saccade tasks were partitioned based on their average spatial accuracy during the two tasks. These partitions (illustrated in Figure 7) formed high error and low error train sets each with $3940$ participants, as well as high error, medium error, and low error test sets each with $245$ participants.
Figure 7: Partitions of users in GazePro based on their average spatial accuracy. The degree value above a partition represents the median spatial accuracy across users in that partition. Note that only the subset of $8615$ users who were recorded for both random saccade tasks are included in these partitions.
EER was highest on the high error test set for both Experiments 16 and 17, which may indicate that poor spatial accuracy reduces authentication performance. Additionally, a model trained on data with high/low error performed worse when evaluated on data with low/high error.
@@
The experiments described above employed $20$ seconds of data for both enrollment and verification. However, it is important to understand how authentication accuracy changes for shorter presentation durations. In this section, we employed presentation durations of $5$, $10$, and $15$ seconds to establish corresponding performance baselines. To do this, instead of computing a centroid embedding over the first four $5$-second chunks of a recording, we used the first one, two, or three chunks. We used the model from Experiment 5 for this analysis.
@@
If we instead used a fixed enrollment duration of $30$ seconds and only varied the verification duration, we observed a similar trend of authentication performance improving with increasing verification duration. We observed that enrolling with $30$ seconds and verifying with $X$ seconds (up to $X=15$) yielded roughly similar results as enrolling and verifying with $X+5$ seconds.
@@
The prior experiments used a random saccade (jumping dot) task for both enrollment and verification. How much, if at all, would performance change by using a smooth pursuit (gliding dot) task instead? We used the same inputs and training/testing sets as Experiment 3 for this analysis, but we trained an ensemble of $4$ models using a $4$-fold split of the training set similar to [27].
@@
For this experiment, we employed the model from Experiment 5 to produce embeddings. We then assessed the permanence (temporal persistence, reliability) of L2-normalized embedding features using the intraclass correlation coefficient (ICC), which has been applied to eye movement traits [10, 11, 33] and is also commonly used in other domains (e.g., [42]).
We found that the median ICC of the $128$ embedding features was $0.94$, with a minimum of $0.90$ and a maximum of $0.98$. Even our lowest ICC of $0.90$ meets all but the strictest of the commonly used thresholds [42], so all $128$ embedding features may be considered highly reliable.
@@ 5 Discussion
(RQ2) We also found that including both the optical and visual axes as inputs to the model yields a substantial improvement in verification accuracy over using the visual axis alone (Experiments 2 and 3). The difference between the visual and optical axes would essentially provide an estimate of a user’s calibration parameters, and this difference alone achieves good verification accuracy (Experiment 4). In a scenario such as continuous authentication where a calibration period could not precede every verification attempt, it may be feasible to employ a cached set of calibration parameters from the enrolled user to estimate the visual axis.
@@
(RQ4) We observed an increase in verification rates as the training population size increased from $250$ to $6000$ users (Experiments 7–14), with the largest performance improvement occurring between $2000$ and $3000$ users. These results suggest that, if there is a point at which performance stops improving when more users are included in the training set, such a point would be beyond $6000$ users. Also, it is possible that training on an even larger population may yield even better verification rates.
(RQ5) We empirically confirmed in practice that, as the test population size increases, verification rates remain stable whereas identification rates decrease (Experiment 15). Eye movement traits are behavioral rather than physical, so they are expected to lack the distinctiveness necessary for large-scale identification. By fitting a curve of the form $a\sqrt{x}+b$ to the observed decline of identification rates and assuming the trend continues, we might expect random-chance identification rates at gallery sizes of about $148000$ identities (see Figure 6).
(RQ6) Regarding the impact of spatial accuracy on authentication performance, we found that poor spatial accuracy reduced authentication performance (Experiments 16–17), in contrast to a study by Raju et al. [33]. Our results also suggest that, for the best performance, the spatial accuracy of the training dataset should be representative of the level of spatial accuracy present in the testing dataset.
(RQ7–8) We observed that, using the model from Experiment 5, $\text{FRR}_{0.002\%}$ is approximately halved for every additional $5$ seconds used for enrollment and verification, starting at $41.9$% with $5$ seconds and improving to $4.1$% with $20$ seconds. Additionally, increasing the enrollment duration improves performance for shorter verification durations. It is important to note, though, that user calibration parameters have still been determined from a full-length calibration process.
@@
(RQ10) The learned embedding features of the model from Experiment 5 all have ICCs at or above $0.9$, indicating that all $128$ embedding features are highly reliable.
@@ 6 Conclusion
In this paper, we investigated authentication performance with a large dataset of gaze recordings collected in VR. Our most important findings indicate that the amount of training data is critical to achieve the authentication performance that is expected from modern wearable devices. Given a very large dataset, which in our case was nearly $10000$ people, it is important to have a longer training time and increased minibatch size. We have highlighted that the best authentication performance can be achieved if binocular estimates of both the optical and visual axes are provided. Crucially, and in support of prior theoretical findings, we found that gaze data alone cannot be employed for (re-)identification purposes if the user pool is sufficiently large. We also discovered the impact of spatial accuracy, an important descriptor of ET signal quality, such that higher signal quality resulted in better authentication performance. In addition, we reinforced the importance of data length on resulting accuracy, specifically that longer enrollment/verification durations improved authentication accuracy. We also found that gaze-based authentication in our study appeared to be stimulus- and task-independent. In summary, gaze authentication can be as accurate as required by the FIDO standard when driven by a state-of-the-art machine learning architecture and a sufficiently large training dataset.
@@ Acknowledgments
We thank Meta for providing us with access to the GazePro dataset and compute resources to perform this important work.

### v2

@@ 3.1 Dataset
Of the $9202$ participants in GazePro, $3673$ self-reported as male, $5376$ as female, and $153$ as neither male nor female. The participants’ ages ranged from $13$ to $88$ years (median=$33$, IQR=$18$). All participants had normal or corrected-to-normal vision. Data collection went through the appropriate IRB approval protocol and received appropriate consent from participants. Additional details about how the dataset was collected and participant characteristics can be found in the supplementary material.
Gaze data in GazePro were collected at $72$ Hz on a VR headset with ET signal quality and other general characteristics equivalent to a Meta Quest Pro [6, 40]. Each recording contains both optical and visual axis estimates from both the left and right eyes. The visual axis is a linear transformation of the optical axis. The parameters of the linear model (essentially, the angle kappa between the optical and visual axes) are fit on a per-user basis via a short calibration procedure.
@@ 3.3 Model evaluation
The first scenario is verification. A presenting user claims to be a specific person (e.g., the owner of a VR headset), their gaze signature is captured, and an accept/reject decision must be made based on the similarity of the gaze signatures captured from the person with the claimed identity at enrollment time (e.g., during account creation) and from the presenting user at verification time. We measure performance in this scenario using equal error rate (EER) and FRR at $1$-in-$50000$ FAR (denoted $\text{FRR}_{0.002\%}$).
The second scenario is identification. A user’s gaze signature is captured, the similarity scores are computed between that user’s gaze signature and all gaze signatures from the population of enrolled participants, and the participant with the highest similarity is determined to be the best match. It is possible to reject a user if the highest similarity score is below some predetermined threshold, in which case it is assumed the user is not part of the enrolled population; but in our case, we evaluated identification rates only on users who were part of the enrolled population. We measure performance in this scenario using Rank-1 identification rate (IR).
@@ 4 Results
We highlight the verification performance under various conditions in Table 1. Unless otherwise noted, results were computed during a random saccade (jumping dot) task with the full training set ($6747$ users) and the full testing set ($2455$ users), using a model trained for $100$ epochs with a minibatch size of $m=256$ samples, and using $20$ seconds of gaze data for both enrollment and verification. Not all $2455$ users in the testing set were recorded for one or both of the tasks used for enrollment/verification, so we had fewer than $2455$ genuine matches. Instead, in this case we obtained $2359$ genuine matches and $5707331$ impostor matches from a matrix of $2385\times 2394$ match scores.
We assigned each experiment a unique ID to enable easier reference. Additional experiments can be found in the supplementary material.
Lohr & Komogortsev [27] used measures of the left eye’s visual axis. As a baseline, we did the same in Experiment 1 and achieved an EER of $8.53$% and a $\text{FRR}_{0.002\%}$ of $99.2$%.
@@
Experiment 2, which used the visual axis from both the left and right eyes, can be compared against our monocular baseline (Experiment 1). This experiment achieved an EER of $3.73$% and a $\text{FRR}_{0.002\%}$ of $93.1$%. These results were substantially better than the monocular baseline.
@@
Experiment 3, which used both the optical and visual axes from both the left and right eyes, can be compared against our visual axis baseline (Experiment 2). This experiment achieved an EER of $0.08$% and a $\text{FRR}_{0.002\%}$ of $17.7$%. These results were substantially better than the visual axis baseline.
@@
We found that increasing the number of training epochs from $100$ to $1000$ yielded a large improvement in performance. Simultaneously increasing the minibatch size from $m=256$ to $m=1024$ further improved performance.
In Experiment 5, we achieved $0.04$% EER and $4.1$% $\text{FRR}_{0.002\%}$ when training for $759$ epochs (terminated early due to time constraints) with a minibatch size of $m=1024$, and when using both the optical and visual axes from both the left and right eyes. These results were substantially better than Experiment 3 which trained for $100$ epochs with a minibatch size of $m=256$. The genuine and impostor similarity distributions and receiver operating characteristic (ROC) curve for this experiment are presented in Figure 3.
Figure 3: Qualitative performance from Experiment 5. The left figure shows histograms of the genuine and impostor similarity score distributions. The right figure shows the ROC curve.
@@
ET signal quality over a large population of people varies for many reasons, including but not limited to eye color, eye shape, presence of corrective lenses, makeup, and device placement on a user’s head [17]. Spatial accuracy, which is defined as the error between a user’s actual gaze and the gaze vector estimated by a gaze estimation pipeline, is one of the most important ET signal quality characteristics. Although a constant error bias in the gaze estimates would not affect measures of velocity, we consider that spatial accuracy is often dependent on gaze angle and might therefore affect features such as saccade curvature.
The subset of $8615$ participants from the full dataset who were recorded for both random saccade tasks were partitioned based on their average spatial accuracy during the two tasks. These partitions (illustrated in Figure 7) formed high error and low error train sets each with $3940$ participants, as well as high error, medium error, and low error test sets each with $245$ participants.
Figure 7: Partitions of a subset of users in GazePro based on their average spatial accuracy. The degree value above a partition represents the median spatial accuracy across users in that partition.
EER was highest on the high error test set for both Experiments 16 and 17, which may indicate that poor spatial accuracy reduces authentication performance. Additionally, a model trained on data with high/low error performed worse when evaluated on data with low/high error. However, we observed a different trend when using the visual axis alone instead of both axes (see the supplementary material).
@@
The prior experiments employed $20$ seconds of data for both enrollment and verification. In this section, we employed presentation durations of $5$, $10$, and $15$ seconds to establish corresponding performance baselines. To do this, instead of aggregating embeddings over the first four $5$-second chunks of a recording, we used the first one, two, or three chunks. We used the model from Experiment 5 for this analysis.
@@
If we instead used a fixed enrollment duration of $30$ seconds, we observed a similar trend of authentication performance improving with increasing verification duration. We observed that enrolling with $30$ seconds and verifying with $X$ seconds (up to $X=15$) yielded roughly similar results as enrolling and verifying with $X+5$ seconds.
@@
The prior experiments used a random saccade (jumping dot) task for enrollment and verification. How much, if at all, would performance change by using a smooth pursuit (gliding dot) task instead? We used the same inputs and training/testing sets as Experiment 3 for this analysis, but we trained an ensemble of $4$ models using a $4$-fold split of the training set similar to [27].
@@
For this experiment, we employed the model from Experiment 5 to produce embeddings. We then assessed the permanence (temporal persistence, reliability) of L2-normalized embedding features using the intraclass correlation coefficient (ICC), which has been applied to eye movement traits [10, 11, 33] and is also commonly used in other domains (e.g., [42]). We found that the median ICC of the $128$ embedding features was $0.94$, with a minimum of $0.90$ and a maximum of $0.98$.
@@ 5 Discussion
(RQ2) We also found that including both the optical and visual axes as inputs to the model yields a substantial improvement in verification accuracy over using the visual axis alone (Experiments 2 and 3). The difference between the visual and optical axes would essentially provide an estimate of a user’s calibration parameters, and this difference alone achieves good verification accuracy (Experiment 4). These results suggest that the angle kappa between the optical and visual axes might be a highly distinctive trait. In a scenario such as continuous authentication where a calibration period could not precede every verification attempt, it may be feasible to employ a cached set of calibration parameters from the enrolled user to estimate the visual axis.
@@
(RQ4) We observed an increase in verification rates as the training population size increased from $250$ to $6000$ users (Experiments 7–14), with the largest performance improvement occurring between $2000$ and $3000$ users. These results suggest that, if there is a point at which performance stops improving when more users are included in the training set, such a point would be beyond $6000$ users, and training on an even larger population may yield even better verification rates.
(RQ5) We empirically confirmed in practice that, as the test population size increases, verification rates remain stable whereas identification rates decrease (Experiment 15). Eye movement traits are behavioral rather than physical, so they are expected to lack the distinctiveness necessary for large-scale identification.
(RQ6) Regarding the impact of spatial accuracy on authentication performance, we found that in some cases poor spatial accuracy reduced authentication performance (Experiments 16–17), in contrast to a study by Raju et al. [33]. Our results also suggest that, for the best performance, the spatial accuracy of the training dataset should be representative of the level of spatial accuracy present in the testing dataset. Further research into the relationship between authentication performance and spatial accuracy is necessary.
(RQ7–8) We observed that, using the model from Experiment 5, $\text{FRR}_{0.002\%}$ is approximately halved for every additional $5$ seconds used for enrollment and verification, starting at $41.9$% with $5$ seconds and improving to $4.1$% with $20$ seconds. Additionally, we observed that increasing the enrollment duration improves performance for shorter verification durations, potentially reducing the time needed for verification attempts at the cost of a longer initial enrollment procedure. It is important to note, though, that user calibration parameters have still been determined from a full-length calibration process.
@@
(RQ10) All 128 learned embedding features of the model from Experiment 5 have ICCs at or above $0.9$, indicating that they are highly reliable according to all but the strictest of the commonly used thresholds [42]. This supports the idea from Friedman et al. [10] that higher ICCs are correlated with higher authentication accuracies.
@@ 6 Conclusion
In this paper, we investigated authentication performance with a large dataset of gaze recordings collected in VR. Our most important findings indicate that the amount of training data is critical to achieve the authentication performance that is expected from modern wearable devices. Given a very large dataset, which in our case was nearly $10000$ people, it is important to have a longer training time and increased minibatch size. We have highlighted that the best authentication performance can be achieved if binocular estimates of both the optical and visual axes are provided. Crucially, and in support of prior theoretical findings, we found that gaze data alone cannot be employed for (re-)identification purposes if the user pool is sufficiently large. We also discovered the impact of spatial accuracy, an important descriptor of ET signal quality, such that higher signal quality resulted in better authentication performance. In addition, we reinforced the importance of data length on resulting accuracy, specifically that longer enrollment/verification durations improved authentication accuracy. We also found that gaze-based authentication in our study appeared to be task-independent. In summary, gaze authentication can be as accurate as required by the FIDO standard when driven by a state-of-the-art deep learning architecture and a sufficiently large training dataset.
@@ Acknowledgments
We thank Meta for providing us with access to the GazePro dataset and the compute resources necessary to perform this important work.